OpenStudy (anonymous):

^3√56x^14 simplify into radical form

OpenStudy (anonymous):

\[\sqrt[3]{56^{14}}\]

OpenStudy (anonymous):

Using the product property of exponents: (a^x)(a^y) = a^(x+y) Do you understand this?

OpenStudy (anonymous):

not really no :/

OpenStudy (anonymous):

\[(a^{x})*(a ^{y})= a ^{x+y}\] How about now?

OpenStudy (anonymous):

If you are multiplying 2 expressions and they share the same base (a), then you can write it as the base with the sum of the exponents

OpenStudy (anonymous):

Kind of, i just dont really understand where the numbers are supposed to go into the formula

OpenStudy (anonymous):

For this question you need to find the greatest multiple of 3 (index) that can go into 11. What is that number?

OpenStudy (anonymous):

Sorry I meant greatest multiple of (3) that can go into 14

OpenStudy (anonymous):

14 is an exponent to 54 though and nothing goes into 14 3 times evenly

OpenStudy (anonymous):

The greatest factor of 3 that can go into 14 is 12. Do you know how I got 12?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

14/3 = 4 with a remainder of 2. Therefore the greatest multiple of 3 that we can get from 14 is 3*(4) = 12 (that does not leave a reaminder)

OpenStudy (anonymous):

i just relized i put the question wrong... ^3√56x^14

OpenStudy (anonymous):

Oh.

OpenStudy (anonymous):

Rewrite: 56 = 8*7 Rewrite x^14 = (x^12)(x^2) \[\sqrt[3]{(2^{3})*(7)*(x^{12})*(x^{2})}\]

OpenStudy (anonymous):

Do you understand how I got this?

OpenStudy (anonymous):

yes